1. Class interval Relative frequenc 0 to < 10 10 to <20 20 to < 30 30 to <40 40 to < 50 0.10 0.35 0.25 0.20 0.10 0.4 0.3 0.2 0.1 0 10 20 30 40 50 What is the value of the median? A. 20 B. 21 C. 22 D. 25 E. 30
Answer #1

Median is middle value

i.e frequency at left of median = frequency at right of median = 0.5

We have (0 to <10) + (10 to <20) = 0.10 + 0.35 = 0.45

To make left side 0.5 we need 0.5-0.45 = 0.05 from (20 to <30) group which is 1/5th of its relative frequency 0.25

Thus, 0.05/0.25 = x / range of group (20 to <30)

Thus, 1/5 = x / 10

x = 2

Thus, median = 20+ 2 = 22

Option C